A soft drink machine outputs a mean of 29 ounces per cup. The machine's output is normally distributed with a standard deviation of 4 ounces. What is the probability of overfilling a 32 ounce cup? Round your answer to four decimal places.
Answer #1

, Given,

= 29, = 4

We convert this to standard normal as

P( X < x) = P( Z < x - / )

P( Z > 32 ) = P( Z > 32 - 29 / 4)

= P( Z > 0.75)

= 0.2266 (Probability calculated from Z table)